OpenStudy (anonymous):

find two numbers whose sum is 24 and whose product is 95

OpenStudy (campbell_st):

so you have x + y = 24 xy = 95 so substitute the 1st equation into the 2nd equation after make y the subject x( 24 - x) = 95 or x^2 - 24x + 95 = 0 solve for x by factoring. hope it helps

OpenStudy (anonymous):

Alternatively .(if you wish to do mentally). factorise 95 as 95 =19*5 and 19+5=24 as here 19 and 5 are primes you may get the answer using hit and trial as above but not in all cases

OpenStudy (anonymous):

solve by factoring: find two numbers whose product is 95 and whose sum is 24 same problem exactly